1. value: 10.00 points The auto repair shop of Quality Motor Company uses standards to control...

1.

value:
10.00 points

The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below:

  Job Standard Hours   Standard Rate   Standard Cost  
  Motor tune-up 1.80         $5.50         $9.90       

The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 170 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups:

  Labor rate variance $ 255 F
  Labor spending variance $ 867 U
Required:
1. Determine the number of actual labor-hours spent on tune-ups during the week.

     

2.

Determine the actual hourly rate of pay for tune-ups last week. (Round your answer to 2 decimal places.)
Answer #1

Labor efficiency variance = 867+255 = 1122 U

a) Labor efficiency variance = (Standard hour-actual hour)Standard rate

-1122 = (170*1.8*5.5-5.5X)

5.5X = 1683+1122

X(actual hour) = 510 Hour

b) Labor rate variance = (Standard rate-actual rate)actual hour

255 = (5.5*510-510X)

510X = 2805-255

X(actual rate) = 5 per hour